OpenStudy (anonymous):

Which is the solution set of the equuation a^2-25=0 A:0,-5 B:0,5 C:-5,5 D:-5,-5

OpenStudy (danjs):

That is a parabola f(x) = a^2 - 25 Setting that to zero f(x) = 0, will give you when y=0, or the x-axis intercepts

OpenStudy (danjs):

since the squared term is +, the thing will open upwards... When a=0, f(a) = -25, that is the y-intercept and the vertex..

OpenStudy (danjs):

 The solution set includes those two points when y=0, where it crosses the 'a' axis

OpenStudy (anonymous):

I dont get it.

OpenStudy (danjs):

you can solve that with algebra, move the 25 over, square root both sides...

OpenStudy (danjs):

recall, \[\sqrt{x^2}= \left| x \right| \]

OpenStudy (danjs):

ill show you for this one.. \[a^2 - 25 = 0\] a^2 = 25 \[\sqrt{a^2}=\sqrt{25}\] \[ \left| a \right| = 5\]

OpenStudy (danjs):

so a can be +5 or -5

OpenStudy (anonymous):

That the answer?

OpenStudy (danjs):

yes, when it asks for the 'solution set'... You just have to find the values for a here. Solving a^2-25=0 for a

OpenStudy (anonymous):

I see

OpenStudy (danjs):

and graphing the thing is a good check... https://www.desmos.com/calculator crosses a axis (y=0) when a is +5 and -5 (use x instead of a to graph it

OpenStudy (danjs):

y=x^2-25

OpenStudy (anonymous):

Isee. thanks